# OpenHuEval: Evaluating Large Language Model on Hungarian Specifics

### v1

@@ B.5 LLM-as-judge
B.6 Comparing the Performance Ranking on WildBench and HuWildBench
We compare the performance rankings of LLM on HuWildBench and WildBench Lin et al. (2024), as shown in Table  7.
@@ Appendix C HuSimpleQA
C.1 Construction pipeline of HuSimpleQA
@@
(1) Obtaining corpora rich in Hungary-specific facts and knowledge: First, we chose the Hungarian Wikipedia website  [^12] https://hu.wikipedia.org/ as the source of corpus material for question construction. We crawled all the entry pages and extracted their content. Next, we used GPT-4o to classify whether the entries were Hungary-specific, with the prompts detailed in Appendix C.2, Figure 17.  [^13] We did not classify all the pages but instead randomly selected pages until we reached 8K Hungary-specific entries, at which point we stopped. We then used GPT to extract key information from the content of these entries suitable for question-answering. The extraction prompts are detailed in Appendix C.2, Figure 18. An example of the extracted key information is shown in Figure 19. As a result, we obtained Hungary-specific key information covering the eight distinct dimensions, totaling 4428 pieces of information.
(2) Generating questions and corresponding answers: We used the GPT-4o model to generate open-ended questions and corresponding answers based on the Hungary-specific key information obtained in the previous step. The prompt used is detailed in Appendix C.3, Figure 20. In this step, we generated a total of 9424 questions based on 4K entries. We then classified the generated questions according to the eight Hungary-specific dimensions outlined in Section 2.1, using GPT-4o with the prompt detailed in Appendix C.3, Figure 21.
(3) Automatic quality checking of questions: To ensure the quality of the questions, we used GPT-4o to check and filter the generated questions. We set the following four criteria, retaining only those questions that met all four standards (the corresponding prompt is detailed in Appendix C.4):
@@
(4) Manual review of question quality: To further ensure the quality of the questions, we hired Hungarian native speakers to manually review the questions. Annotators checked whether the questions met the four criteria mentioned in Step 3. During the annotation process, each question was assigned to two annotators, who received the questions but not the answers. A question was considered valid and retained only if both annotators agreed that it met all four criteria and that the provided answer matched the original reference answer. Detailed procedures are provided in Appendix C.4. After these four steps, we obtained a total of 1293 questions, with their distribution across the eight Hungary-specific dimensions shown in Table 16.
C.2 Obtaining corpora rich in Hungary-specific facts and knowledge
@@
C.3 Generating questions and corresponding answers
@@
C.4 Automatic quality checking of questions
@@
C.5 Manual review of question quality
@@ C.7 LLM-as-judge
[…]
@@ 5 QwQ GPT-4o-mini(↑3)
C.8 Comparing the Performance Ranking on HuSimpleQA and SimpleQA
We compare the performance rankings of LLM on HuSimpleQA and SimpleQA Wei et al. (2024), as shown in Table  9.
Appendix D HuProverbRea
D.1 Construction pipeline of HuProverb
Figure 12: Construction of HuProverbRea.
D.2 More examples of HuProverbRea
@@ 5 Deepseek-R1 Deepseek-V3 (↓2) o1-mini (↑2) Deepseek-V3 (↓2)
D.3 Comparing the Performance Ranking on HuProverbRea and MAPS
We compare the performance rankings of LLM on HuProverbRea and MAPS Liu et al. (2024c), as shown in Table  10 and  11.
Appendix E HuMatchingFIB and HuStandardFIB
E.1 Construction of HuMatchingFIB and HuStandardFIB
@@
E.2 More examples of HuMatchingFIB and HuStandardFIB
@@
E.3 Metric and Judge of HuMatchingFIB and HuStandardFIB
@@
HuStandardFIB questions are designed with open-ended reference answers to accommodate variations in part of speech and semantics. We employ a many-to-one fuzzy matching mechanism. Fuzzy matching is a technique that calculates the similarity between strings, allowing for flexibility in matching by considering variations such as typos, synonyms, or different word orders. In this context, the model’s answer is compared against a set of possible reference answers (where multiple correct answers may exist for a single question or blank). If the similarity score between the model’s answer and any of the reference answers exceeds a predefined threshold, the answer is considered correct. This approach is particularly suitable for evaluating open-ended questions where exact matches are often infeasible due to the variability in acceptable responses. The annotator information involved in all tasks of OpenHuEval can be found in Appendix G.
Appendix F Analyzing LRM’s thinking process on OpenHuEval
F.1 Analyzing LRM’s thinking process on HuSimpleQA
@@
F.2 Analyzing LRM’s thinking process on HuMatchingFIB
F.2.1 Splitting LRM’s thinking process into segments and classifying these segments
@@
F.2.2 Tagging the reasoning segments along the dimensions
@@
Appendix G Information of the Annotators

### v2

@@ D.5 LLM-as-judge
D.6 Comparing the Performance Ranking on WildBench and HuWildBench
We compare the performance rankings of LLM on HuWildBench and WildBench Lin et al. (2024), as shown in Table  9.
@@ Appendix E HuSimpleQA
E.1 Construction pipeline of HuSimpleQA
@@
(1) Obtaining corpora rich in Hungary-specific facts and knowledge: First, we chose the Hungarian Wikipedia website  [^12] https://hu.wikipedia.org/ as the source of corpus material for question construction. We crawled all the entry pages and extracted their content. Next, we used GPT-4o to classify whether the entries were Hungary-specific, with the prompts detailed in Appendix E.2, Figure 17.  [^13] We did not classify all the pages but instead randomly selected pages until we reached 8K Hungary-specific entries, at which point we stopped. We then used GPT to extract key information from the content of these entries suitable for question-answering. The extraction prompts are detailed in Appendix E.2, Figure 18. An example of the extracted key information is shown in Figure 19. As a result, we obtained Hungary-specific key information covering the eight distinct dimensions, totaling 4428 pieces of information.
(2) Generating questions and corresponding answers: We used the GPT-4o model to generate open-ended questions and corresponding answers based on the Hungary-specific key information obtained in the previous step. The prompt used is detailed in Appendix E.3, Figure 20. In this step, we generated a total of 9424 questions based on 4K entries. We then classified the generated questions according to the eight Hungary-specific dimensions outlined in Section 2.1, using GPT-4o with the prompt detailed in Appendix E.3, Figure 21.
(3) Automatic quality checking of questions: To ensure the quality of the questions, we used GPT-4o to check and filter the generated questions. We set the following four criteria, retaining only those questions that met all four standards (the corresponding prompt is detailed in Appendix E.4):
@@
(4) Manual review of question quality: To further ensure the quality of the questions, we hired Hungarian native speakers to manually review the questions. Annotators checked whether the questions met the four criteria mentioned in Step 3. During the annotation process, each question was assigned to two annotators, who received the questions but not the answers. A question was considered valid and retained only if both annotators agreed that it met all four criteria and that the provided answer matched the original reference answer. Detailed procedures are provided in Appendix E.4. After these four steps, we obtained a total of 1293 questions, with their distribution across the eight Hungary-specific dimensions shown in Table 22.
E.2 Obtaining corpora rich in Hungary-specific facts and knowledge
@@
E.3 Generating questions and corresponding answers
@@
E.4 Automatic quality checking of questions
@@
E.5 Manual review of question quality
@@ E.7 LLM-as-judge
$\text{F-Score}=\frac{2}{\frac{c+i}{c}+\frac{c+i+n}{c}}=\frac{2c}{2c+2i+n}$ (2)
@@
  Model Judge Correct Ratio
  Deepseek-R1 0.98
Deepseek-V3 1.00
Qwen2.5-Instruct-7B 0.98
Qwen2.5-Instruct-72B 0.98
Llama-3.1-Instruct-8B 1.00
Llama-3.1-Instruct-70B 1.00
GPT-4o 0.99
GPT-4o-mini 1.00
o1-mini 1.00
QwQ 0.98
Table 10: Accuracy of LLM-as-Judge for HuSimpleQA.
To verify the effectiveness of LLM-as-judge, we randomly selected 100 samples from each of the 10 evaluated models for inspection, with accuracy exceeding 98% for all models, as shown in Table 10.
@@ 5 QwQ GPT-4o-mini(↑3)
E.8 Comparing the Performance Ranking on HuSimpleQA and SimpleQA
We compare the performance rankings of LLM on HuSimpleQA and SimpleQA Wei et al. (2024), as shown in Table  12.
E.9 Results of Translating HuSimpleQA to English
  Model Original HuSimpleQA Translated to EN
  GPT-4o 50.3 35.96 (↓14.34)
GPT-4o-mini 24.52 19.72 (↓4.80)
QwQ 9.17 12.84 (↑3.67)
Deepseek-R1 34.58 32.02 (↓2.56)
Deepseek-V3 32.71 29.08 (↓3.63)
Llama-3.1-70B 36.36 23.82 (↓12.54)
Llama-3.1-8B 14.9 14.08 (↓0.82)
o1-mini 16.24 14.31 (↓1.93)
Qwen2.5-72B 15.05 8.89 (↓6.16)
Qwen2.5-7B 5.29 15.47 (↑10.18)
Table 13: Results of translating HuSimpleQA to English.
As shown in Table 5, the current leading LLMs perform poorly on HuSimpleQA. To further analyze and confirm whether these issues stem from the models’ lack of knowledge of the Hungarian language or Hungarian-specific cultural aspects, we translated HuSimpleQA into English and reevaluated the models. The results, presented in Table 13, show that after translation, 2 out of 10 models improved, while 8 experienced a decline in performance, indicating that translation generally worsens performance. This suggests that LLMs have more difficulty understanding Hungarian cultural nuances than the language itself. These findings are consistent with the conclusions in
E.10 Human Performance on SimpleQA and HuSimpleQA
  Model HuSimpleQA-abs HuSimpleQA-rel SimpleQA-abs SimpleQA-rel
GPT-4o 55.33 0.85 38.0 1.12
GPT-4o-mini 24.87 0.38 8.7 0.26
Deepseek-V3 33.8 0.52 24.9 0.73
QwQ 14.82 0.23 21.0 0.62
Deepseek_R1 35.75 0.55 30.1 0.89
Qwen2.5-7B-Instruct 5.92 0.09 5.3 0.16
Qwen2.5-72B-Instruct 16.07 0.25 12.2 0.36
Llama-3.1-8B-Instruct 15.66 0.24 12.6 0.37
Llama-3.1-70B-Instruct 37.34 0.58 29.3 0.86
o1-mini 26.88 0.41 11.3 0.33
AVG 26.65 0.40 19.34 0.57
Table 14: The results of HuSimpleQA and SimpleQA, abs is the absolute performance of CGA metric, rel is the performance compared to human.
We hire human annotators to complete the tasks in the SimpleQA and HuSimpleQA datasets in order to assess human performance on these datasets. HuSimpleQA was annotated by Hungarian native speakers, while SimpleQA was handled by English native speakers. We used two annotators per dataset, averaging their performance metrics to determine human performance. Annotators were instructed NOT to use external tools like search engines or GPT. For HuSimpleQA and SimpleQA, they were told to answer "I don’t know" if uncertain about a question.
The results for HuSimpleQA and SimpleQA are shown in Table 14. The average absolute performance (CGA) of the ten models is 26.65 vs 19.34, highlighting significant room for improvement in LLM performance on both datasets. Regarding average relative performance to humans, the scores are 0.40 vs 0.57, indicating that LLMs perform much worse on HuSimpleQA compared to SimpleQA. Despite better absolute performance on HuSimpleQA, LLMs are significantly weaker in simple factual ability in Hungarian than in English.
@@ 5 Deepseek-R1 Deepseek-V3 (↓2) o1-mini (↑2) Deepseek-V3 (↓2)
Appendix F HuProverbRea
F.1 Construction pipeline of HuProverb
Figure 12: Construction of HuProverbRea.
F.2 More examples of HuProverbRea
  Model HuProverbRea-abs HuProverbRea-rel MAPs(en)-abs MAPs(en)-rel
Deepseek-V3 92.51 1.15 96.95 1.02
Deepseek_R1 91.72 1.14 98.48 1.03
GPT-4o 95.51 1.19 97.97 1.03
GPT-40-mini 92.16 1.15 97.97 1.03
Llama-3.1-70B-Instruct 93.83 1.17 98.48 1.03
Llama-3.1-8B-Instruct 73.48 0.91 92.39 0.97
o1-mini 87.67 1.09 92.89 0.97
Qwen2.5-72B-Instruct 90.22 1.12 98.22 1.03
Qwen2.5-7B-Instruct 67.05 0.83 94.16 0.99
QwQ 84.23 1.05 67.51 0.71
AVG 86.84 1.08 93.50 0.98
Table 17: The results of HuProverbRea and MAPs(en), abs is absolute performance, rel is the performance compared to human.
F.3 Comparing the Performance Ranking on HuProverbRea and MAPS
We compare the performance rankings of LLM on HuProverbRea and MAPS Liu et al. (2024c), as shown in Table  15 and  16.
F.4 Human Performance on HuProverbRea and MAP(en)
We hire human annotators to complete the tasks in the HuProverbRea and MAP(en) datasets in order to assess human performance on these datasets. HuProverbRea was annotated by Hungarian native speakers, while MAP(en) was handled by English native speakers. We used two annotators per dataset, averaging their performance metrics to determine human performance. Annotators were instructed NOT to use external tools like search engines or GPT. The accuracy for HuProverbrea is 80.44%. The accuracy for MAP(en) is 95.31%.
Furthermore, we calculated the LLMs’ relative performance to human, as shown in Table  17.
The average absolute performance of the ten models is 0.8684 vs 0.9350, indicating that LLMs perform worse on HuProverbRea compared to English proverb reasoning, suggesting room for improvement in Hungarian proverb tasks.
For average relative performance to humans, the scores are 1.08 vs 0.98. This indicates that while LLMs slightly underperform compared to humans on the MAP(en) dataset, they outperform humans on the HuProverbRea dataset, highlighting their strengths in Hungarian language-related knowledge and reasoning.
Appendix G HuMatchingFIB and HuStandardFIB
G.1 Construction of HuMatchingFIB and HuStandardFIB
@@
G.2 More examples of HuMatchingFIB and HuStandardFIB
@@
G.3 Metric and Judge of HuMatchingFIB and HuStandardFIB
@@
HuStandardFIB questions are designed with open-ended reference answers to accommodate variations in part of speech and semantics. We employ a many-to-one fuzzy matching mechanism. Fuzzy matching is a technique that calculates the similarity between strings, allowing for flexibility in matching by considering variations such as typos, synonyms, or different word orders. In this context, the model’s answer is compared against a set of possible reference answers (where multiple correct answers may exist for a single question or blank). If the similarity score between the model’s answer and any of the reference answers exceeds a predefined threshold, the answer is considered correct. This approach is particularly suitable for evaluating open-ended questions where exact matches are often infeasible due to the variability in acceptable responses. The annotator information involved in all tasks of OpenHuEval can be found in Appendix I.
Appendix H Analyzing LRM’s thinking process on OpenHuEval
H.1 Analyzing LRM’s thinking process on HuSimpleQA
@@
H.2 Analyzing LRM’s thinking process on HuMatchingFIB
H.2.1 Splitting LRM’s thinking process into segments and classifying these segments
@@
H.2.2 Tagging the reasoning segments along the dimensions
@@
Appendix I Information of the Annotators